How do I solve this equation?
12c + 8.25(55 – c) = 585

Answer by ewatrrr(24785)


Hi
solvIng
12c + 8.25(55 – c) = 585
3.75c = 585 - 8.25*55
3.75c = 131.25
c = 35
CHECKING our Answer***
12*35 + 8.25*20 = 585